OpenStudy (anonymous):

What is the simplified form of 1 over x minus 2 over x squared plus x

OpenStudy (anonymous):

\[\frac{ 1 }{ x }-\frac{ 2 }{ x }^2+x\]

OpenStudy (anonymous):

Is that your question?

OpenStudy (anonymous):

Hello? @hamza123

OpenStudy (anonymous):

yeah

OpenStudy (anonymous):

okay

OpenStudy (anonymous):

i think the answer is 1-x/x( x+1)

OpenStudy (anonymous):

\[\frac{ x-3 }{ x } \]

OpenStudy (anonymous):

i think that's the answer :)

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

that is wrong

OpenStudy (anonymous):

okay sorry